OpenStudy (anonymous):

Tyrone has a total of 50 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have?

OpenStudy (anonymous):

Tyrone has a total of 50 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have? A. 12 B. 38 C. 36 D. 14 Its d

OpenStudy (anonymous):

C + D = 50 (2 + 3D) + D = 50 -> 4D = 48 => D = 12 DVDs => C = 2 + 36 = 38 CDs